10. A small child is eating dinner on a flat table that is 1 meter tall. When asked to eat his vegetables, the child pushes his plate away from him at a speed of 4 m/s. Due to friction, the plate decelerates at a constant rate of 1 m/s2. After 1 second, the plate falls off the edge of the table.


a.     How far did the plate travel over the 1 second?

 

 

 

  

 

b.     What is the speed of the plate at the instant it falls off the table?

 

 

 

 

 

c.     How much time will pass between the instant the plate leaves the table and the instant it reaches the ground?

 

 

 

 

 

d.     At what speed will the plate strike the ground?

 

 


 

 

e.     At what angle will the plate strike the ground?

 




1
Expert's answer
2021-01-22T11:28:44-0500

a)


"s=4(1)-(0.5)(1)(1)^2=3.5\ m"

b)


"v=4-(1)(1)=3\frac{m}{s}"

c)


"h=0.5gt^2\\1=0.5(9.8)t^2\\t=0.45\ s"

d)


"V=\sqrt{3^2+(9.8\cdot0.45)^2}=5.3\frac{m}{s}"

e)


"\theta=\arctan{\frac{3}{9.8\cdot 0.45}}=34\degree"